Equipment was acquired at the beginning of the year at a cost of $78,660. The equipment was depreciated using the straight-line method based on an estimated useful life of six years and an estimated residual value of $7,920. Required:

(a) What was the depreciation expense for the first year?

(b) Assuming the equipment was sold at the end of the second year for $59,486, determine the gain or loss on sale of the equipment.

(c) Journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

a.

Depreciation expense = $78,660 - 7,920 / 6 = $11,790

b.

Book value at the end of second year = $78,660 - 23,580 (11,790*2) = $55,080

Gain on sale of equipment = $59,486 - 55,080 = $4,406

c.

General Journal Debit Credit
Cash $59,486
Accumulated depreciation ($11,790*2) 23,580
Gain on sale of equipment $4,406
Equipment 78,660